A particle of mass m and charge q is moving along the x-axis with
speed u when it passes between two charged parallel plates. The plates generate a uniform
electric field E_0 j in the region 0 ≤ x ≤ b and no field elsewhere. Find the angle through which the particle is deflected by its passage between the plates.
1
Expert's answer
2019-04-17T11:13:23-0400

Force, applied to the particle:


"F = q E_0"

Force parallel to the y-axis.

Acceleration of the particle:


"a = a_y = \frac{F}{m} = \frac{qE_0}{m}"

Time of passing this region with electric field:


"t = \frac{b}{u}"

Speed of the particle in y-direction


"u_y = a_y \cdot t = \frac{ q E_0 b}{mu}"

Therefore, angle of deflection:


"\tan \theta = \frac{u_y}{u}=\frac{ q E_0 b}{mu^2}"